log7 + log (n - 2) = log 6n
Answer by mukhopadhyay(490)

log7 + log (n - 2) = log 6n
=>log[7(n-2)] = log 6n
=> 7(n-2) = 6n
=> 7n-14 = 6n
=> n = 14